Find two consecutive even integers such that the smaller added to three times the larger gives a sum of 38.


1
Expert's answer
2021-03-11T06:10:31-0500

Let x represent the smaller integer and y the bigger integer. Since they are consecutive even integers, we have that

"y= x +2"

Also, interpreting the given data, we have that

"x +3y = 38\\ x +3(x+2) = 38\\ x + 3x + 6 = 38\\ 4x + 6 = 38\\ 4x = 32 \\ x = 8 \\ \therefore y = 10"

So the two consecutive even integers are 8 and 10.